OpenStudy (anonymous):

how do i solvem-4n=8 for m

OpenStudy (anonymous):

To solve for "m", all you have to do is have "m" by itself on one side.

OpenStudy (anonymous):

divide both sides by -4

Directrix (directrix):

m - 4n = 8 for m m -4n + 4n = 8 + 4n m = 8 + 4n

OpenStudy (anonymous):

I don't understand exactly what you mean by "how" considering you don't know the value for n, you can put it in equation form which is to move the 4n to the right hand side

OpenStudy (anonymous):

thank you what about f+4 and under that g and that =6 for f